### v1

# SolarGPT-QA: A Domain-Adaptive Large Language Model for Educational Question Answering in Space Weather and Heliophysics

Abstract
Solar activity, including solar flares, coronal mass ejections (CMEs), and geomagnetic storms, can significantly impact satellites, aviation, power grids, data centers, and space missions. Extreme solar events can cause substantial economic damage if not predicted in advance, highlighting the importance of accurate forecasting and effective education in space science. Although large language models (LLMs) perform well on general tasks, they often lack domain-specific knowledge and pedagogical capability to clearly explain complex space science concepts.
We introduce SolarGPT-QA, a question answering system based on a domain-adapted large language model built on the LLaMA-3 base model. The model is trained using scientific literature and large-scale question-answer data generated with GPT-4 and refined using Grok-3 in a student-friendly storytelling style. Human pairwise evaluations show that SolarGPT-QA outperforms general-purpose models in zero-shot settings and achieves competitive performance compared to instruction-tuned models for educational explanations in space weather and heliophysics. A small pilot student comprehension study further suggests improved clarity and accessibility of the generated explanations. Ablation experiments indicate that combining domain-adaptive pretraining with pedagogical fine-tuning is important for balancing scientific accuracy and educational effectiveness. This work represents an initial step toward a broader SolarGPT framework for space science education and forecasting.
I Introduction
Heliophysics provides a system-level understanding of the Sun and its influence throughout the heliosphere [34]. Within this broader heliophysical framework, space weather science investigates the complex and dynamic interactions between the Sun, the solar wind, interplanetary space, and planetary environments, and examines how solar activity influences Earth’s space environment and technological systems [45]. Solar phenomena such as solar flares (sudden bursts of electromagnetic radiation from the Sun) [45, 44], coronal mass ejections (CMEs) (large eruptions of magnetized plasma from the solar corona) [54], and solar energetic particles (SEPs) (high-energy charged particles accelerated by flares and CMEs) [42, 2] can trigger geomagnetic storms [19], which disrupt satellite operations, data centers, communication systems, and power grids, and pose significant radiation risks to astronauts and high-altitude aviation. For instance, the ”Carrington Event” in 1859 [33] was the most intense geomagnetic storm on record triggered by an extreme solar flare and CME, the Quebec blackout in 1989 [14] showed the vulnerability of modern technological system to space weather events.
Such disruptions have resulted in economic losses of billions of dollars and, in extreme cases, pose risks to human safety and critical infrastructure [5, 2]. As modern society becomes increasingly dependent on space-based technologies—such as large satellite constellations (e.g., Starlink) [38] and as private space enterprises including SpaceX and Blue Origin pursue future space-based data centers and exploration missions [52], understanding space weather has become a critical scientific, technological, and educational priority, beginning as early as the high-school level.
Large language models have become powerful tools for understanding and explaining complex information, with growing impact across many scientific fields [3]. They can answer questions, summarize research, and support learning. However, their use in space weather and heliophysics is still limited, largely because these areas rely on highly specialized literature and concepts that are rarely covered in general training data. One key challenge is most space weather knowledge lives in technical papers and mission reports, which means general models often miss key terms, cause–effect links, or physical limits. Another issue is teaching space science particularly to K–12 students or early undergraduates requires clear language, analogies, and storytelling approach. General models are not built to balance accuracy that is tailored to the appropriate age.
To fill this gap, we introduce SolarGPT-QA, a domain-adapted model designed for educational question answering in space weather and heliophysics for K–12 learners. The model is trained on a carefully curated corpus of high-impact, peer-reviewed research articles published from 2010 onward, aligned with the Solar Dynamics Observatory (SDO) era of modern flare and CME research. These sources are further refined for relevance to solar eruptions and used to generate a large collection of synthetic question–answer pairs. Rather than targeting expert-level analysis, SolarGPT-QA focuses on helping students grasp core concepts through clear, accessible explanations to questions such as how solar flares form or why the Sun can disrupt technology on Earth.
Figure 1: Overview of the SolarGPT-QA system illustrating its educational question–answering framework for space weather and heliophysics.
II Related Work
II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [18]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [28, 37, 20]. Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [53, 57], but their operational use is limited by computational cost and incomplete boundary conditions [56].
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [21] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters, employing decision trees and sequence models to model temporal dependencies [31].
Progress in this area has been supported by the development of benchmark datasets. The SWAN-SF multivariate time series dataset, integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [1]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [16, 22].
Many efforts have used deep learning on solar imagery, including magnetograms and extreme ultraviolet observations, to capture spatial precursors to flares and CMEs, while multimodal fusion of time series and images further enhances forecasting accuracy [36, 58, 23]. Despite these advances, most operational systems predict flares, CMEs, and SEPs independently and provide limited explanatory context, functioning largely as black boxes [13, 39, 40]. These limitations motivate unified, multimodal, and explainable forecasting frameworks that build on prior advances in time series modeling, dataset development, and interpretability.
II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve best performance across wide range of the tasks and model scales [6, 12, 50, 51, 9, 8, 7]. There has been subsequent work and exploring for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [41, 60, 49, 43, 59]. DAPT and finetuning are common techniques to improve the LLM for specific dataset or specific domain. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [26, 24, 17]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [55, 35, 47, 11].
Beyond general-purpose models, large language models have increasingly been adapted to specialized scientific domains, including medicine, biology, Earth science, and environmental sciences. In the medical domain, Med-PaLM [46] demonstrates that large-scale language models can encode clinical knowledge and perform competitively on medical question answering through instruction tuning and expert evaluation. Generalist medical foundation models, such as those proposed by Moor et al. [32], extend this approach by enabling a single model to support a wide range of clinical reasoning tasks. MedGPT [25] leverages electronic health records and named entity recognition to predict future medical events, while BioGPT [30] is pre-trained on large-scale biomedical literature to improve biomedical text generation and knowledge extraction.
Geneformer [48] is a life science domain-adapted language model that applies transformer-based language modeling to single-cell transcriptomic data, enabling transfer learning for gene network inference and cellular state prediction. Similarly, large-scale protein language models demonstrate that language modeling techniques can predict atomic-level protein structures directly from amino acid sequences [29].
More recently, scientific LLMs have been introduced for Earth, ocean, and geoscience applications. GeoGPT [15] represents one of the first large language models explicitly designed for geoscience tasks. OceanGPT [4] is trained on marine and oceanographic literature, demonstrating superior performance and preliminary embodied intelligence for diverse ocean science and engineering tasks compared to general-purpose LLMs. CLLMate [27] focuses on climate-related applications, integrating climate science literature and data-driven knowledge to support climate analysis, interpretation, and decision support. Tele-Knowledge Pretraining [10] incorporates structured technical knowledge into language models for fault diagnosis and engineering analysis.
Despite these advances, most existing scientific LLMs primarily target expert-level analysis and professional decision-making, often producing technically dense and jargon-heavy outputs. In contrast, our work focuses on domain-adapted language modeling for educational scientific question answering, emphasizing conceptual clarity, accessibility, and age-appropriate explanations. Rather than supporting expert users, SolarGPT-QA is explicitly designed to facilitate learning and understanding in complex scientific domains such as space weather and heliophysics.
III SolarGPT-QA
SolarGPT-QA is a domain-adapted large language model designed specifically for educational question answering in space weather and heliophysics, with an emphasis on K–12–friendly explanations. The methodology consists of four main stages: (1) domain corpus construction and preprocessing, (2) domain-adaptive pretraining (DAPT) using LoRA, (3) educational question–answer fine-tuning, and (4) inference-time constraints to promote clarity and reduce hallucinations. An overview of the system architecture is shown in Fig. 2. Let $\mathcal{D}_{H}$ denote a corpus of heliophysics literature and $\mathcal{D}_{QA}$ denote a supervised question–answer dataset designed for K–12 learners. Our objective is to learn a conditional language model
$p_{\theta}(y\mid x),$ (1)
where $x$ is a heliophysics question and $y$ is an age-appropriate explanatory answer that preserves scientific correctness while prioritizing conceptual clarity.
Figure 2: System architecture and workflow of SolarGPT-QA, showing the data sources, domain adaptation, and question–answer generation pipeline.
III-A Domain Corpus Construction and Normalization
We curate a heliophysics corpus $\mathcal{D}_{H}=\{d_{i}\}_{i=1}^{N}$ from peer-reviewed literature published between 2010 and 2024, aligned with the operational era of the Solar Dynamics Observatory (SDO). Each document $d_{i}$ is extracted from PDF format and transformed using a deterministic preprocessing function:
$\tilde{d}_{i}=f_{\text{clean}}(d_{i}),$ (2)
where $f_{\text{clean}}$ removes references, citations, section headers, figure captions, URLs, and formatting artifacts, followed by Unicode normalization.
The cleaned corpus $\mathcal{T}=\{\tilde{d}_{i}\}$ is tokenized into sequences
$\mathbf{x}=(x_{1},\dots,x_{T}),\quad x_{t}\in\mathcal{V},$ (3)
where $\mathcal{V}$ denotes the tokenizer vocabulary of the base language model.
III-B Domain-Adaptive Pretraining with Low-Rank Adaptation
We initialize SolarGPT-QA from a pretrained causal language model, Meta-LLaMA-3-8B, and adapt it to the heliophysics domain using domain-adaptive pretraining (DAPT) combined with Low-Rank Adaptation (LoRA) for parameter efficiency. As defined in Section III-A, the heliophysics corpus is processed as
$DH\rightarrow\mathcal{T}$
DAPT minimizes the autoregressive negative log-likelihood over the tokenized corpus $\mathcal{T}$:
$\mathcal{L}_{\text{DAPT}}(\theta)=\mathbb{E}_{x\sim\mathcal{T}}\left[\sum_{t=1}^{T}-\log p_{\theta}(x_{t}\mid x_{<t})\right]$ (4)
This objective biases the model toward heliophysics-specific terminology, causal structures, and semantic patterns associated with solar flares, coronal mass ejections (CMEs), and solar energetic particles (SEPs).
To enable parameter-efficient adaptation, we employ Low-Rank Adaptation (LoRA). For a transformer weight matrix
$W\in\mathbb{R}^{d\times k},$
LoRA applies a low-rank update
$W\rightarrow W^{\prime}=W+\Delta W,$ (5)
where
$\Delta W=BA.$ (6)
The overall training workflow is summarized in Algorithm 1. Only the low-rank matrices $A$ and $B$ are trainable, while the base model parameters remain frozen. This significantly reduces the number of optimized parameters and enables efficient domain adaptation without sacrificing the general linguistic capabilities of the pretrained model. Specific LoRA configuration values are provided in Section IV.
III-C Fine-Tuning on QA dataset
We define a supervised dataset
$\mathcal{D}_{QA}=\{(q_{i},a_{i})\}_{i=1}^{M},$ (7)
where each answer $a_{i}$ is rewritten to emphasize analogies, simplified causal explanations, and age-appropriate language.
Fine-tuning minimizes the conditional likelihood:
$\mathcal{L}_{QA}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D}_{QA}}\left[-\log p_{\theta}(a\mid q)\right].$ (8)
During this stage, all base model parameters remain frozen and only LoRA parameters are updated. Inputs are formatted explicitly as:
Q: <question>
A: <answer>
to align the model with educational discourse patterns.
Algorithm 1 Training Procedure for SolarGPT-QA
1:Require: Heliophysics corpus $\mathcal{D}_{H}$, QA dataset $\mathcal{D}_{QA}$, pretrained model $p_{\theta_{0}}$
2: Clean and normalize $\mathcal{D}_{H}\rightarrow\mathcal{T}$
3: Initialize LoRA adapters on attention layers
4:Domain-Adaptive Pretraining:
5:for each batch $\mathbf{x}\in\mathcal{T}$ do
6:  Update LoRA parameters using $\mathcal{L}_{\text{DAPT}}$
7:end for
8:Educational Fine-Tuning:
9:for each $(q,a)\in\mathcal{D}_{QA}$ do
10:  Format input as Q: q A: a
11:  Update LoRA parameters using $\mathcal{L}_{QA}$
12:end for
13:return Trained SolarGPT-QA model
The overall training objective is:
$\mathcal{L}=\mathcal{L}_{\text{DAPT}}+\lambda\mathcal{L}_{QA},$ (9)
where $\lambda$ controls alignment toward educational clarity.
Unlike prior scientific language models optimized for expert-level analysis, SolarGPT-QA is explicitly designed for educational deployment. By combining domain-adaptive pretraining, parameter-efficient optimization, and constrained decoding, the model bridges expert heliophysics knowledge and accessible K–12 instruction.
IV Experiment
IV-A Training Configuration and Hyperparameters
All experiments were conducted using LoRA-based fine-tuning while keeping the base Meta-LLaMA-3-8B parameters frozen. LoRA adapters were applied to the query and value projection matrices of the attention layers. The rank of the low-rank adapters was set to $r=16$, with a scaling factor $\alpha=32$ and dropout probability $p=0.05$.
During both domain-adaptive pretraining and fine-tuning, optimization was performed using the AdamW optimizer. At inference time, text generation employed stochastic decoding with a temperature of $\tau=0.7$, top-$k=50$, and nucleus sampling with top-$p=0.9$. Generation was terminated at sentence boundaries using a custom stopping criterion, and post-processing filtered incomplete sentence fragments to ensure syntactic completeness and readability. All of our code used in this experiment is available on GitHub [^1] https://github.com/hlabcsusu/SolarGPT.
IV-B Dataset
The datasets used in this study consist of two components: (1) a heliophysics domain corpus for domain-adaptive pretraining (DAPT) and (2) a question–answer (QA) dataset for educational fine-tuning and evaluation.
IV-B1 Domain Corpus for DAPT
The heliophysics corpus is curated from peer-reviewed research articles published between 2010 and 2024, aligned with the operational era of the Solar Dynamics Observatory (SDO). Articles were selected from high-impact journals with an h5-index greater than 40 to ensure scientific rigor and relevance. These venues include The Astrophysical Journal (ApJ), The Astrophysical Journal Supplement Series (ApJS), Astronomy & Astrophysics (A&A), Space Weather, and Frontiers in Astronomy and Space Sciences, among others.
Papers were filtered using eruption-related keyword co-occurrence such as solar flare, coronal mass ejection (CME), and solar energetic particles (SEP). Figure 3 shows the distribution of eruption-related research articles across five leading heliophysics journals between 2010 and 2024, stratified by individual and combined event types. This analysis identifies 1,570 relevant papers from these venues alone, and extending the selection to all qualifying journals yields an estimated 5,000–7,000 research articles used for domain adaptation.
Figure 3: Combination of Eruption Event Types as Paper Keywords
All documents were extracted from PDF format and cleaned by removing references, citations, figure captions, URLs, and formatting artifacts. Corpus-level statistics confirm strong coverage of eruption-related concepts. In particular, the word cloud shown in Fig. 4 highlights the dominance of key heliophysics terms such as solar, CME, flare, SEP, magnetic field, particle, and shock, indicating that the corpus effectively captures the semantic structure and terminology of solar eruptive phenomena.
IV-B2 QA Dataset Construction
The supervised QA dataset was constructed by generating question–answer pairs from the curated heliophysics literature using large language models, followed by pedagogical rewriting to emphasize clarity, storytelling, and age-appropriate explanations suitable for K–12 learners. Questions focus on conceptual understanding rather than mathematical derivations, targeting physical intuition, cause–effect relationships, and real-world impacts of space weather phenomena.
IV-B3 Train–Test Split and Difficulty Levels
The training set consists of QA pairs derived from the same corpus used during domain-adaptive pretraining. In contrast, the test set is constructed from QA pairs extracted from different PDF documents that are not included in the training corpus, ensuring document-level separation and reducing information leakage. To support a fine-grained educational evaluation, the test set is further divided into three levels of difficulty:
  • Easy: Basic factual and definitional questions, such as identifying solar phenomena or simple cause–effect relationships, suitable for elementary and middle school learners.
  • Medium: Conceptual reasoning questions that require explanation of physical processes or interactions (e.g., how CMEs interact with Earth’s magnetosphere), aligned with middle to early high school understanding.
  • Hard: Integrative questions that require synthesizing multiple concepts or explaining observational consequences (e.g., relationships between flares, CMEs, and SEPs), suitable for advanced high school students.
This tiered evaluation design enables systematic assessment of SolarGPT-QA across increasing levels of conceptual complexity while maintaining consistent scientific grounding.
Figure 4: WordCloud of the data corpus used for DAPT
IV-C Baseline
We evaluate SolarGPT-QA against several strong general-purpose large language models that are widely used for scientific question answering and education. The baselines include ChatGPT, Grok, Claude, DeepSeek, and Gemini. All baseline models are evaluated under a zero-shot setting using identical prompts and decoding parameters, and without retrieval augmentation, to ensure a fair and controlled comparison.
In addition, we include an instruction-tuned variant of ChatGPT, where the model is prompt-tuned using a small number of representative examples (few-shot setting), in order to compare SolarGPT-QA against a highly optimized general-purpose model with explicit pedagogical alignment.
IV-D Evaluation Method
We use win rate as the primary evaluation metric based on human pairwise comparison. For each test question, a human evaluator compares SolarGPT-QA’s response with a baseline model’s response and assigns one of three outcomes: Win, Tie, or Loss. The win rate is defined as the proportion of questions where SolarGPT-QA is judged to outperform the baseline. Results are reported by aggregating win, tie, and loss rates across all questions and separately across difficulty levels (Easy, Medium, Hard).
IV-E Results
We evaluate model performance using human pairwise comparison to assess response quality in educational question answering for space weather and heliophysics. Figure 5 presents the overall win-rate comparison across multiple baselines.
ChatGPT (instruction-tuned, 10-shot) achieves the highest win rate among all evaluated models, indicating that human evaluators consistently prefer its responses. This result highlights the substantial benefit of instruction tuning and few-shot prompting for improving clarity and response relevance. SolarGPT-QA performs competitively and outperforms all other zero-shot and general-purpose baselines, but remains inferior to ChatGPT when enhanced with instructional alignment. ChatGPT (zero-shot) and Grok demonstrate moderate performance, while Claude, DeepSeek, and Gemini obtain lower win rates, underscoring the challenges faced by general-purpose models in specialized educational contexts.
Figure 5: Human evaluation win-rate comparison across models.
To further assess practical educational effectiveness, we conduct a small pilot student comprehension study involving 15 students. Each student was asked to read answers generated by each model and report whether they understood the response. Figure 6 shows the number of students who reported understanding the answers produced by each model.
ChatGPT (instruction-tuned, 10-shot) achieves the highest level of student comprehension, with all 15 students reporting understanding its responses. SolarGPT-QA follows closely, with 12 students indicating comprehension. ChatGPT (zero-shot) and Grok demonstrate moderate comprehension levels, while Claude, Gemini, and DeepSeek show lower student understanding overall. These results further suggests that instructional alignment and pedagogical optimization play a crucial role in educational question answering, beyond domain specialization alone.
Figure 6: Student comprehension comparison across models.
IV-F Ablation Study
We conduct an ablation study to analyze the impact of domain-adaptive pretraining (DAPT) and educational fine-tuning in SolarGPT-QA. We compare four model variants: the base LLaMA-3 model, LLaMA-3 with DAPT only, LLaMA-3 with QA fine-tuning only, and the full SolarGPT-QA model. DAPT significantly improves scientific grounding and correct use of heliophysics terminology but often produces explanations that are too technical for K–12 learners. QA fine-tuning alone improves readability and storytelling but lacks sufficient domain depth, leading to shallow or incomplete explanations. Table I shows SolarGPT-QA achieves the highest win rate (75%) and student comprehension (80%), suggesting that combining domain-adaptive pretraining with educational fine-tuning yields the strongest performance.
TABLE I: Ablation Study Results for SolarGPT-QA
Model Variant Win Rate (%) Student Understanding (%)
Base LLaMA-3-8B 28 20 (3/15)
LLaMA-3 + DAPT 54 40 (6/15)
LLaMA-3 + QA Fine-Tuning 48 67 (10/15)
SolarGPT-QA 75 80 (12/15)
The results indicate that SolarGPT-QA benefits significantly from domain-adaptive pretraining and fine-tuning. While ChatGPT and Grok often generate factually correct responses, they frequently use advanced terminology or omit explanatory details, which reduces their educational effectiveness.
V Discussion and Conclusion
V-A Discussion
The results presented in this work demonstrate that domain adaptation alone is insufficient for educational effectiveness, and that pedagogical alignment is a critical and often overlooked dimension when deploying large language models for science education. While general-purpose LLMs such as ChatGPT and Grok exhibit strong linguistic fluency and factual recall, our findings show that they frequently fail to deliver explanations that are accessible, coherent, and age-appropriate for K–12 learners in the context of space weather and heliophysics.
In the human pairwise evaluation, instruction-tuned ChatGPT (10-shot) achieves the highest overall win rate, indicating that evaluators consistently favor responses that are explicitly aligned with instructional intent and teaching style. Despite this, SolarGPT-QA performs competitively and outperforms all zero-shot general-purpose baselines, achieving a 75% overall win rate. This result suggests that domain-adaptive pretraining combined with educational fine-tuning can substantially narrow the gap between zero-shot domain models and heavily instruction-tuned general LLMs, without relying on in-context examples at inference time.
The student comprehension study further refines this interpretation. While instruction-tuned ChatGPT achieves perfect comprehension among participants, SolarGPT-QA follows closely, with a large majority of students reporting clear understanding of its explanations. In contrast, comprehension levels for zero-shot general-purpose models remain noticeably lower. These results indicate that SolarGPT-QA succeeds in translating complex heliophysics concepts into accessible explanations, even if it does not surpass the strongest instruction-aligned baseline.
The ablation study provides critical insight into why this performance emerges. Domain-adaptive pretraining (DAPT) alone improves scientific grounding but tends to preserve expert-level language, while QA fine-tuning alone improves readability at the expense of domain depth. The full SolarGPT-QA model achieves the strongest balance between these objectives, producing explanations that are both scientifically grounded and pedagogically accessible. Importantly, SolarGPT-QA reaches this balance without external prompting strategies, highlighting the effectiveness of integrating educational objectives directly into the training process.
These findings have broader implications beyond heliophysics. They suggest that many existing scientific LLMs—while impressive for expert-level analysis—are poorly suited for educational deployment without explicit adaptation. As AI systems increasingly enter classrooms, museums, and public outreach platforms, models that fail to account for learner expertise risk reinforcing misunderstanding or disengagement. SolarGPT-QA provides preliminary evidence that targeted domain adaptation combined with educational fine-tuning can effectively bridge this gap.
V-B Future work
The present study represents an initial, proof-of-concept exploration of domain-adapted large language models for educational question answering in space weather and heliophysics. Several important extensions are planned as part of the broader SolarGPT research agenda.
First, future work will focus on developing SolarGPT-Core, a large-scale domain-adapted foundation model trained on an expanded corpus of peer-reviewed heliophysics literature. Scaling to larger architectures and longer context windows is expected to improve scientific depth, reasoning consistency, and coverage of rare but high-impact solar events.
Second, we plan to introduce SolarGPT-PredEx, a multimodal prediction-and-explanation model that integrates time-series measurements and solar imagery with language modeling. Unlike the current SolarGPT-QA system, SolarGPT-PredEx will reason directly over observational data to produce both event predictions and data-grounded natural language explanations, enabling more transparent and interpretable space weather forecasting.
Third, future versions of SolarGPT-QA will incorporate retrieval-augmented generation (RAG), curriculum-aware prompting, and adaptive difficulty control to support classroom deployment and personalized learning. Larger and more diverse human evaluation studies—particularly involving teachers and students across multiple grade levels—will be conducted to rigorously assess educational effectiveness.
Finally, all components of the SolarGPT ecosystem will be integrated into a unified research and outreach platform that bridges prediction, explanation, and education, supporting both scientific discovery and public understanding of heliophysics. The current work establishes the feasibility of this vision while clearly delineating its preliminary scope.
V-C Conclusion
In this work, we introduced SolarGPT-QA, a domain-adapted large language model designed for educational question answering in space weather and heliophysics. By combining domain-adaptive pretraining on peer-reviewed heliophysics literature with pedagogically aligned question–answer fine-tuning, SolarGPT-QA aims to bridge expert scientific knowledge and accessible K–12 instruction.
Human pairwise evaluation results show that while instruction-tuned ChatGPT (10-shot) achieves the strongest overall performance, SolarGPT-QA outperforms all zero-shot general-purpose baselines and suggests strong student comprehension and explanatory clarity. While these findings are preliminary and based on limited sample sizes, they provide initial evidence that integrating domain adaptation with educational fine-tuning can enhance the clarity and accessibility of scientific explanations. Ablation experiments further demonstrate that neither domain-adaptive pretraining nor educational fine-tuning alone is sufficient; rather, their combination is important for balancing scientific accuracy with pedagogical effectiveness.
Beyond its direct use, this work shows a clear method for building educational science language models. It highlights that good AI learning tools should include teaching goals, not just subject knowledge. SolarGPT-QA is a first step toward a larger SolarGPT system that brings together prediction, explanation, and education in heliophysics.
VI Acknowledgements
Shah Muhammad Hamdi is supported by the GEO directorate under NSF awards #2301397 and #2530946. Soukaina Filali Boubrahimi is supported by GEO Directorate under NSF awards #2204363, #2240022, and #2530946.